# On the Volatility of Shapley-Based Contribution Metrics in Federated Learning

### v3

@@ Abstract
Federated learning (FL) is a collaborative and privacy-preserving Machine Learning paradigm, allowing the development of robust models without the need to centralize sensitive data. A critical challenge in FL lies in fairly and accurately allocating contributions from diverse participants. Inaccurate allocation can undermine trust, lead to unfair compensation, and thus participants may lack the incentive to join or actively contribute to the federation. Various remuneration strategies have been proposed to date, including auction-based approaches and Shapley-value-based methods, the latter offering a means to quantify the contribution of each participant. However, little to no work has studied the stability of these contribution evaluation methods. In this paper, we evaluate participant contributions in federated learning using gradient-based model reconstruction techniques with Shapley values and compare the round-based contributions to a classic data contribution measurement scheme. We provide an extensive analysis of the discrepancies of Shapley values across a set of aggregation strategies, and examine them on an overall and a per-client level. We show that, between different aggregation techniques, Shapley values lead to unstable reward allocations among participants. Our analysis spans various data heterogeneity distributions, including independent and identically distributed (IID) and non-IID scenarios.
@@ I Introduction
Federated Learning (FL), a collaborative method of training Machine Learning models without the involved entities having to share their datasets, has gained a lot of traction in recent years as a prevalent form of collaborative learning [1]. The growing focus on privacy preservation, driven by the enactment of regulations such as the General Data Protection Regulation (GDPR), emphasizes the importance of FL in addressing data privacy concerns.
@@
In addition to the performance of the model, the evaluation and fair distribution of rewards to the participants in a federation is of the utmost importance, particularly in commercial settings where financial incentives might be needed for the participation of the client [6]. Each participant would expect to be fairly compensated for their participation, regardless of the underlying aggregation strategy used.
In this work, we use cumulative round-based Shapley values as percentages to represent the contribution of each client. Shapley values, a concept from cooperative game theory, are employed to quantify the contribution of each participant in a federation by evaluating their marginal impact on the model’s performance in each communication round. This approach is a popular contribution evaluation method [7] and reflects a practical industry use case where, in every round, the central server assesses the contribution of each client based on their local model updates. Over time, the server aggregates these round-based contributions to determine the overall contribution of each participant, enabling fair reward distribution. We analyze, across an array of popular aggregation strategies, the overall, as well as per-client, reliability of these generated contribution values.
@@ II-A Federated Learning
A basic Federated Learning architecture consists of a group of participants, or clients, who each own a private local dataset, along with a central server or aggregator. Before training starts, they agree upon a model architecture and hyperparameters such as learning rates, the aggregation strategy, and local epoch numbers.
@@ II-B Aggregation strategies
  • Finally, Krum [11], similarly to the two previous methods, is a modified aggregation to counter adversarial clients. It excludes client updates by excluding those that are in general too far away from the other clients.
@@ IV-A Experimental Setup
An open-source repository containing an implementation of Shapley value calculations in FL, used in our experiments is available at https://github.com/arnogeimer/flwr-contributions.
@@ IV-A3 Strategies
We employ a collection of 8 different aggregation strategies, which have all been widely adopted across their respective application domains. While aggregation strategies that do not have the same objective function as FedAvg have been proposed [28], we do not include any of them in our analysis. Performance metrics of strategies, whether related to model accuracy or defense against attacks, are not the primary focus of this paper. We therefore employ out-of-the-box hyper-parameters in all strategies.
@@ IV-A4 Technical specifications
As we used 4 different datasets and three Dirichlet splitting values, the experiments included 12 different use cases. With 70 seeds and 3 epoch values $e\in(2,5,10)$ per use case, and a total of 8 aggregation strategies, our experiments encompass more than 20.000 full Federated Learning runs.
@@ IV-B1 Round-based aggregation
We will call the contribution of the client $k$ the normalized weighted sum of their per-round Shapley values $\phi_{k}^{t}$. A participant’s total contribution is thus represented on a percentage scale, allowing for better comparisons between runs: Although the numerical values of the Shapley values may vary, the proportions remain comparable. Shapley values are aggregated using an inverse linear factor to the rounds, up to a maximum contribution halting round $R$. For example, at $R=10$, the Shapley values are added using weight $\frac{10}{10}$ in round $1$, $\frac{9}{10}$ in round 2, etc:
@@ IV-B2 Ground truth
Determining which aggregation strategy produces the values closest to a ground truth is of the utmost importance. To this end, we establish the ground-truth contribution relative to the sizes of the participants’ datasets. Since we do not operate in an adversarial environment, where actors may lie about their dataset statistics, and the distribution of classes across different splittings is non-heterogeneous, we deem the size of splits to be a reliable contribution assessment for clients. This method has previously been used in the literature[31][32]. The correctness of a contribution is calculated using the squared Euclidean distance to the ground truth, in line with previous studies. To better understand our results, we establish a lower bound using an equal payout contribution evaluation approach.
@@ VI Contribution instabilities between aggregation strategies
In addition, the figure contains $\alpha$-specific histograms. We observe that there is no value that performs significantly better than the others; in fact, the error distribution is stable between different values. This illustrates that no matter how heterogeneous the data, the problem persists. This is undesirable in a non-malicious federation: A client receiving almost half of the total reward more depending on the aggregation technique is not in the interest of anyone involved.
@@
The scale of the problem is shown in Figure 3, demonstrating the instability between different strategies. Ideally, all histograms would show no variance, with all differences close to zero. We show that this is not the case: Although the variance between strategies changes, there always exist substantially different contribution allocations. While there are clearly some strategies which agree with others on most contributions, especially FedAvg, we show that there are still discrepancies of up to $30\%$ of the total reward when comparing with other aggregations. In summary, we have shown that, even if the overall contribution lies close to the ground truth, different aggregation strategies will lead to differing contributions per client. As described above, this can cause substantial problems when establishing federations.
@@ VII Discussion
Therefore when configuring federation parameters in a real-world setting, it is crucial to ensure that any transition to an alternative aggregation strategy is accompanied by a comprehensive stability study such as the one presented in this work. Through such a study, federations can ensure that the new aggregation strategy not only performs well in terms of model accuracy but also provides contribution stability, fostering a sustainable and trustworthy continuation of the federation.
On a more general context, we advocate for a reorientation in the design and evaluation of federated learning contribution systems. Beyond traditional metrics like accuracy, future research and development efforts should include fairness and robustness of contributions in their studies, to foster a sustainable and trustworthy federated learning ecosystem.

### v4

@@ Abstract
Federated learning (FL) is a collaborative and privacy-preserving Machine Learning paradigm, allowing the development of robust models without the need to centralize sensitive data. A critical challenge in FL lies in fairly and accurately allocating contributions from diverse participants. Inaccurate allocation can undermine trust, lead to unfair compensation, and thus participants may lack the incentive to join or actively contribute to the federation. Various remuneration strategies have been proposed to date, including auction-based approaches and Shapley-value-based methods, the latter offering a means to quantify the contribution of each participant. However, little to no work has studied the stability of these contribution evaluation methods. In this paper, we evaluate participant contributions in federated learning using gradient-based model reconstruction techniques with Shapley values and compare the round-based contributions to a classic data contribution measurement scheme. We provide an extensive analysis of the discrepancies of Shapley values across a set of aggregation strategies and examine them on an overall and a per-client level. We show that, between different aggregation techniques, Shapley values lead to unstable reward allocations among participants. Our analysis spans various data heterogeneity distributions, including independent and identically distributed (IID) and non-IID scenarios.
@@ I Introduction
Federated Learning (FL), a collaborative method for training Machine Learning models without the entities involved having to share their datasets, has gained a lot of traction in recent years as a prevalent form of collaborative learning [1]. The growing focus on privacy preservation, driven by the enactment of regulations such as the General Data Protection Regulation (GDPR), emphasizes the importance of FL in addressing data privacy concerns.
@@
In addition to the performance of the model, the evaluation and fair distribution of rewards to participants in a federation is of utmost importance, particularly in commercial settings where financial incentives might be needed for the participation of the client [6]. Each participant would expect to be fairly compensated for their participation, regardless of the underlying aggregation strategy used.
In this work, we use cumulative round-based Shapley values as percentages to represent the contribution of each client. Shapley values, a concept from cooperative game theory, are employed to quantify the contribution of each participant in a federation by evaluating their marginal impact on the model’s performance in each communication round. This approach is a popular contribution evaluation method [7] and reflects a practical industry use case where, in each round, the central server assesses the contribution of each client based on their local model updates. Over time, the server aggregates these round-based contributions to determine the overall contribution of each participant, allowing a fair reward distribution. We analyze, across an array of popular aggregation strategies, the overall, as well as per-client, reliability of these generated contribution values.
@@ II-A Federated Learning
A basic Federated Learning architecture consists of a group of participants, or clients, who each own a private local dataset, along with a central server or aggregator. Before training starts, they agree on the model architecture and hyperparameters such as learning rates, the aggregation strategy, and local epoch numbers.
@@ II-B Aggregation strategies
  • Finally, Krum [11], similarly to the two previous methods, is a modified aggregation to counter adversarial clients. It excludes client updates that are too far away from the other clients.
@@ IV-A Experimental Setup
An open source repository containing an implementation of Shapley value calculations in FL, used in our experiments, is available at https://github.com/arnogeimer/flwr-contributions.
@@ IV-A3 Strategies
We employ a collection of 8 different aggregation strategies, which have all been widely adopted across their respective application domains. Although aggregation strategies that do not have the same objective function as FedAvg have been proposed [28], we do not include any of them in our analysis. Performance metrics of strategies, whether related to model accuracy or defense against attacks, are not the primary focus of this paper. We therefore employ out-of-the-box hyperparameters in all strategies.
@@ IV-A4 Technical specifications
As we used 4 different datasets and three Dirichlet splitting values, the experiments included 12 different use cases. With 70 seeds and 3 epoch values $e\in(2,5,10)$ per use case and a total of 8 aggregation strategies, our experiments encompass more than 20.000 full Federated Learning runs.
@@ IV-B1 Round-based aggregation
We will call the contribution of the client $k$ the normalized weighted sum of their per-round Shapley values $\phi_{k}^{t}$. The total contribution of a participant is thus represented on a percentage scale, allowing better comparisons between runs. Although the numerical values of the Shapley values may vary, the proportions remain comparable. Shapley values are aggregated using an inverse linear factor for rounds, up to a maximum contribution halting round $R$. For example, at $R=10$, the Shapley values are added using weight $\frac{10}{10}$ in round $1$, $\frac{9}{10}$ in round 2, etc:
@@ IV-B2 Ground truth
Determining which aggregation strategy produces the values closest to a ground truth is of the utmost importance. To this end, we establish the ground-truth contribution relative to the sizes of the participants’ datasets. Since we do not operate in an adversarial environment, where actors may lie about their dataset statistics, and the distribution of classes across different splittings is nonheterogeneous, we deem the size of splits to be a reliable contribution assessment for clients. This method has previously been used in the literature[31][32]. The correctness of a contribution is calculated using the squared Euclidean distance to the ground truth, in line with previous studies. To better understand our results, we establish a lower bound using an equal payout contribution evaluation approach.
@@ VI Contribution instabilities between aggregation strategies
In addition, the figure contains $\alpha$-specific histograms. We observe that there is no value that performs significantly better than the others; in fact, the error distribution is stable between different values. This illustrates that no matter how heterogeneous the data, the problem persists. This is undesirable in a nonmalicious federation: A client receiving almost half of the total reward more depending on the aggregation technique is not in the interest of anyone involved.
@@
The scale of the problem is demonstrated in Figure 3, which shows the instability between different strategies. Ideally, all histograms would show no variance, with all differences close to zero. We show that this is not the case: Although the variance between strategies changes, there always exist substantially different contribution allocations. Although there are clearly some strategies that agree with others in most contributions, especially FedAvg, we show that there are still discrepancies of up to $30\%$ of the total reward when compared to other aggregations. In summary, we have shown that, even if the overall contribution lies close to the ground truth, different aggregation strategies will lead to differing contributions per client. As described above, this can cause substantial problems when establishing federations.
@@ VII Discussion
Therefore, when configuring federation parameters in a real-world setting, it is crucial to ensure that any transition to an alternative aggregation strategy is accompanied by a comprehensive stability study such as the one presented in this work. Through such a study, federations can ensure that the new aggregation strategy not only performs well in terms of model accuracy but also provides contribution stability, fostering a sustainable and trustworthy continuation of the federation.
In a more general context, we advocate for a reorientation in the design and evaluation of federated learning contribution systems. Beyond traditional metrics like accuracy, future research and development efforts should include the fairness and robustness of contributions in their studies to foster a sustainable and trustworthy federated learning ecosystem.